The Photolysis of Acetylene Halides Reaction with Oxygen

\underline{Mei-Lee Hwang}, Deaprtment of Chemical Engineering, Kaohsiung Polytechnic Institute, Taiwan, 84008; Yu-Ping Kuo, Department of Applied Chemistry, Chia Nan Junior College of Pharmacy, Taiwan.

The photolysis of chloro- and bromo-acetylene reactions at low temperature with Hg lamp are carried out. There is no new absorption band was observed after in-situ photolysis of acetylene halides in Ar matrix. With the presence of oxygen, several new bands were appeared. The new peaks at 2138 and 2342 cm-1 were assigned as CO and CO2, respectively. Further work are essential for the assignment of the other new bands.